OpenStudy (anonymous):

a – 1, a, a + 1, a + 2, a + 3, __________ a) a + 4 b) a – 4 c) a + 5 d) a – 3

OpenStudy (anonymous):

It's a) a+4

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Let see: \(\large (a – 1) + 1=a \\\large a+1=a+1 \\\large (a+1)+1=a+2 \\\large (a+2)+1=a+3\\\large so\ (a+3)+1= \color{blue}{a+4}\)

OpenStudy (anonymous):

ohh.. ok.. thanx :)

OpenStudy (anonymous):

\(\huge \color{green}{WELCOME}\)